Express the following as the sum of two odd primes. 
44

Grade:12

4 Answers

Harshit Singh
3 years ago
Dear Student

3 + 41 = 44

Thanks
Jitendra Kumar
3 years ago
Dear Student,
44 = 3 + 41
Also it can be written as
44 = 7 + 37
44 = 13 +31
I hope youbt is cleared
Thanks
Ram Kushwah
3 years ago
Sum of two odd primes is always an even number
But all odd numbers are not prime 
Here are the possible combination of prime numbers 
whose sum is 44
 
3+41=44
7+37=44
13+31=44
Thanks 
pl aprove
 
 
Shraddha Singh
2 years ago
Dear student,
The possible combinations of odd primes to get 44 as sum are
 
3 + 41= 44
7 + 37 = 44
13 + 31= 44
 
Hope it helps!
Thank you